# AuthPrint: Fingerprinting Generative Models Against Malicious Model Providers

### v1

@@ Results
Another contributor to the performance gap is the generative process itself. GANs leave subtle artifacts or statistical footprints that AuthPrint may implicitly exploit (Marra et al. 2019). Diffusion models, by contrast, suppress such artifacts through iterative denoising. As shown in Figure 9 in Appendix B, reducing denoising steps substantially lowers detection error in SD—suggesting that shallower sampling may preserve more model-specific traces.
Scaling Up Training Enhances Detection
@@
As shown in Figure 4, we evaluated reconstructors with 32M, 187M, and 674M parameters on SD models, holding all other variables constant. The smallest (32M) failed to reject negative samples, indicating insufficient capacity for the fingerprinting task. In contrast, performance improved substantially with larger reconstructors, and the 674M model achieved near-perfect detection error on SD 1.1.
We also varied reconstructor’s training set size (128k, 320k, 512k samples) and observed vanishing detection error with more data. These results show that scaling both reconstructor capacity and training volume significantly improves performance. In our most demanding setting (674M reconstructor, 512k samples), training took $\sim$240 GPU hours on an NVIDIA H200—costly for academic clusters but feasible in industrial or governmental settings. Importantly, training is only needed once per target model.
@@
Boosting Performance with Ensemble Detection
To reduce FPR, AuthPrint can be naturally extended with an ensemble detection scheme, where multiple reconstructors are independently trained using different fingerprint index vectors. At verification, each image is evaluated by all reconstructor-index pairs, and the final detection error is defined as the worst-case MSE across the ensemble, i.e., $e(x^{\prime})=\max_{i\in[N]}\text{MSE}(R_{\phi_{i}}(x^{\prime}),(x^{\prime})_{s_{i}})$, where each $R_{\phi_{i}}$ is trained on $s_{i}$, and $N$ is the ensemble size. This conservative strategy boosts sensitivity: while a negative sample may yield low error on one reconstructor-index pair, it is unlikely to do so across all, thereby reducing false acceptances.
@@
We evaluate this scheme on SD models (Figure 5) using identical hyperparameters but distinct index sets per reconstructor. Detection error decreases with ensemble size, with diminishing returns suggesting convergence. The ensemble design is flexible and can be extended to heterogeneous configurations with varying fingerprint lengths, architectures, or training setups. We leave exploration of such to future work.
Detection of Target Model Modifications
@@
Figure 6 shows the relationship between detection error (FPR) and image quality (FID) under all perturbations. AuthPrint achieves effective trade-offs under pruning and downsampling. For example, moderate pruning (62.5%) increases FID modestly to 14.69 but drives FPR to near-zero. Similarly, downsampling from $\times256$ to $\times224$ sharply reduces FPR despite only slight increases in FID. Quantization, by contrast, appears more robust to detection: high precision loss (INT4) renders high FID (>> 200) and leads to 0 FPR as expected, but moderate quantization (INT6) results in high FPR despite quality degradation (FID $\approx$ 53). This indicates that detection is less effective for quantization than for pruning or downsampling under the same fingerprint configuration. These results indicate that while AuthPrint is effective at detecting distribution shifts, its sensitivity varies by perturbation type.
Prompt Specificity Enhances Detection for SD
@@
This raises a key question for AuthPrint: How does prompt specificity affect detection performance? We evaluate detection error under five prompt regimes, ordered by increasing specificity (see Appendix A for prompt details):
@@
Figure 7: Effect of prompt specificity on detection performance for SD. We compare five regimes: I. Diverse prompts from DiffusionDB; II. Category-specific “People” prompts from PartiPrompts; III. A focused subset of II (see Appendix A); IV. A single generic prompt (“a high quality photo”); V. A single specific prompt (Japanese café prompt).
@@
These results reflect an intuitive pattern: diverse prompts lead to a broad, overlapping output space, making it harder for the detector to distinguish models. Specific prompts constrain the output distribution, making model differences more detectable. Importantly, strong performance only for specific prompts does not limit practical applicability. Many real-world applications—e.g., profile portraits, product photos, architectural renderings, or avatars—rely on consistent, repeatable prompts, where our approach can be effective.
Evaluation of Adaptive Adversarial Attacks
Evasion Attacks.
A common strategy for bypassing detectors is evasion, such as attacks based on Projected Gradient Descent (PGD) (Madry et al. 2017). In our setting, evasion translates to forgery, where an adversary perturbs an image sampled from a negative model to evade detection. We evaluate this threat in the model substitution scenario, using a reconstructor trained with a fingerprint length of 32 on a target StyleGAN2 model (FFHQ, 70k, ADA), against the four negative models shown in Figure 2 (left) that yield exactly 0% FPR@95%TPR. The adversary has full access to both target and negative models. It trains a classifier to distinguish their outputs and uses it to guide a PGD attack on the detector. PGD parameters are in Appendix A.
@@
    Baseline: A ResNet-18 classifier trained to distinguish target vs. negative outputs without using any fingerprint.
@@
    Yu-2019: A state-of-the-art model fingerprinting approach where fingerprints are public and not output-dependent (Yu, Davis, and Fritz 2019).
@@
    AuthPrint: Our proposed method using a reconstructor trained on secret, output-dependent fingerprints.
With AuthPrint, the attacker operates in a regime where the detector details are kept secret. To mount the attack, they assume the reconstructor has the same architecture as the baseline classifier, train a surrogate detector, and use its gradients to guide PGD. For each detector, we report the attack success rate, average number of steps to success, and LPIPS perceptual distance of the forged images.
Results show that AuthPrint is immune to this attack: the attacker fails to generate even a single successful forgery. As shown in Table 1, Appendix C, both baseline detectors are defeated by PGD, with a 100% attack success rate—i.e., all non-authentic images are misclassified as authentic. Moreover, the attacks preserve perceptual quality, with LPIPS consistently <<0.01. While the Yu-2019 detector requiress slightly more steps to evade, both are ultimately ineffective. In contrast, AuthPrint achieves a 0% attack success rate. The attacker’s inability to access the secret pixel indices or the reconstructor prevents them from training a viable surrogate. This secrecy is essential: if the attacker had white-box access to the reconstructor, or knowledge of the fingerprint index vector, a PGD attack would likely succeed.
Fingerprint Recovery Attacks.
Beyond evasion-based forgery, an adversary may attempt to recover the fingerprint indices $s$. We decompose this attack into two steps: first, recovering the fingerprint set $S:={\mathrm{set}(s)}$; second, searching the ordered vector $s$ from $S$. The order is crucial—without it, the adversary cannot replicate the behavior of $R$.
Step 1: Recover the set $S$. To recover the set $S$, one can probe the detector through pixel manipulation: starting from an authentic image $x$, the attacker manipulates a subset $M\subset[d]$ of $m$ pixels (e.g.setting them to 0) to produce a modified image $x^{\prime}$, and observes whether it passes or is rejected.
The detector computes $D(x)=\mathrm{MSE}[R(x),x_{s}]$, which stays low if two conditions hold: (1) $R(x)$ is stable, which happens when $m$ is small (see Figure 11, Appendix D), and (2) the manipulated pixels do not significantly overlap with the fingerprint set $S$. In other words, the overlap $|M\cap S|$ must be large enough to corrupt $x_{s}$ for the MSE to increase and lead to rejection. We refer to this required overlap as rejection threshold $k$ (see Appendix D for how $k$ is estimated).
This rejection behavior reduces Step 1 to the non-adaptive threshold group testing (NATGT) problem, where the fingerprint $S$ plays the role of the hidden “defective” set, and each attacker query corresponds to a test set $M_{i}$. The detector returns positive (rejection) iff $|M_{i}\cap S|\geq k$. The attacker’s goal is to recover $S$ using as few queries as possible.
Equation 1 in Appendix D reflects recent theoretical estimation for the query complexity $T$ from NATGT applied to our problem. For instance, with $d=3\cdot 256^{2}$, $\ell=32$, and $k=2$, it yields $T=\mathcal{O}(10^{5})$ queries, which is impractical for detectors under oversight, e.g., enforcing rate limits or access controls to prevent abuse. $T$ exhibits approximately exponential growth with respect to the length ratio $k/l$ (see Figure 10 in Appendix D). Importantly, this estimation assumes the attacker can query with arbitrary $m$, whereas our threat model restricts to small $m$. As a result, this serves as a lower bound on recovery complexity. Tight estimates under limited queries are out of the scope of this work.
An ensemble of detectors can further boost robustness: accepting if any detector accepts reduces false negatives and makes recovery harder. This differentiates from the earlier max-MSE ensemble strategy, which prioritized minimizing false positives. The choice reflects verifier goals and trade-offs between robustness and performance.
Step 2: Recover the vector $s$. If the attacker learns $S$ but not its order in $s$, they cannot train an effective surrogate reconstructor $\hat{R}$ because a different order would result in fingerprint mismatches on out-of-distribution samples. To recover $s$ from $S$ requires searching over all permutations, $O(l!)$. For example, when $l=32$, this corresponds to roughly $O(10^{35})$. Hence, Step 2 dominates for longer fingerprints, while Step 1 dominates for shorter fingerprints.
Discussion and Conclusion
With AI regulations coming into force, model providers in sensitive domains may be subject to compliance audits, thereby creating a demand for accountability and transparency. To address this, the academic community has proposed mechanisms to attribute generated outputs to the specific model that produced them. However, these approaches face challenges for a scalable and practical deployment.
We have tackled this problem for image generation models by adapting model fingerprinting techniques and investigating their robustness in adversarial settings. Our proposed approach allows a verifier to detect whether a given image was sampled from a model of quality lower than that of a reference model. This scheme does not require specialized hardware or modifications to the reference model, and is compatible with modern generative algorithms, offering a more practical alternative to existing approaches.
We evaluate the scheme on two popular generative models—StyleGAN2 and Stable Diffusion—and observe low detection error for images generated by models trained with sub-optimal configurations, older versions from the same model family, and compressed model variants. Furthermore, the error steadily decreases with more training data and larger reconstructor models, suggesting that verifiers can trade computational resources for improved performance. We also show that the scheme is robust to two types of adversarial forgery attacks: gradient-based evasion and fingerprint recovery.
Although the scheme does not offer the formal guarantees of cryptographic methods, it explores a complementary direction by leveraging statistical differences in output distributions. In real-world scenarios, such evidence may suffice to raise concerns or prompt further investigation. To mitigate attacks, we recommend the verifier follows standard operational security practices to protect the fingerprint indices and reconstructor, enforces rate limits, and deploys API authentication and reputation mechanisms.
We view this work as a step toward practical, scalable verification tools for generative models, especially while cryptographic and hardware-based solutions are still maturing. Promising future directions include ensemble methods, techniques for amplifying detection confidence, and approaches to improve performance on conditioned models.
Appendix A Experimental Methodology
We evaluate AuthPrint on two classes of pre-trained image generators: StyleGAN2 and Stable Diffusion (SD).
Pre-trained Generative Models.
StyleGAN2 models are sourced from the official StyleGAN2-ADA repository by NVLabs [^2] https://github.com/NVlabs/stylegan2-ada-pytorch and pre-trained on two datasets: FFHQ and LSUN-Cat. All models generate fixed-size $256\times 256$ outputs. We experiment with several variants trained on different datasets and/or augmentation strategies.
For model substitution evaluations:
  • FFHQ models: The target model is trained on 70k images with ADA (Adaptive Discriminator Augmentation). The four negative models are: (1) 1k data with ADA, (2) 30k data with ADA, (3) 70k data with BCR (Balanced Consistency Regularization), and (4) 70k data with no augmentation.
  • LSUN-Cat models: The target model is trained on 100k images with ADA. The four negative models are: (1) 1k data with ADA, (2) 30k data with ADA, (3) 100k data with BCR, and (4) 100k data with no augmentation.
In each case, negative models differ from the target model either in data volume or augmentation method.
SD 2.1 and SD 1.5–1.1 models are obtained from HuggingFace [^3] https://huggingface.co, and by default generate $1024\times 1024$ images across all experiments.
Fingerprint Reconstructor.
The AuthPrint reconstructor is a convolutional neural network that takes a full generated image as input and predicts the pixel values at a secret subset of spatial pixel locations (channel-aware) of this image. It is trained to recover these values solely from the full image itself.
We evaluate three reconstructor sizes: 32M, 187M, and 674M parameters. Unless otherwise noted, the 32M model is used for StyleGAN2 experiments and the 674M model for SD experiments, with all sizes included in the ablation study on reconstructor capacity.
Across all sizes, the architecture follows a common design: a deep convolutional encoder with progressively increasing channel widths and stride-2 downsampling, implemented via stacked convolutional blocks using $\times4$ kernels, batch normalization, and LeakyReLU activations. The encoder is followed by a global average pooling layer and a multi-layer MLP head. The model takes an image as input and outputs a vector representation of the fingerprint.
Training Data for Reconstructors.
For StyleGAN2, training images are generated by sampling latent vectors $z\sim\mathcal{N}(0,I)$. For SD, unless otherwise noted, training uses a single fixed prompt—referred to as the Japanese café prompt—described below and illustrated in Figure 8:
“A photorealistic photo for a Japanese café named NOVA CAFE, with the name written clearly on a street sign, a storefront banner, and a coffee cup. The scene is set at night with neon lighting, rain-slick streets reflecting the glow, and people walking by in motion blur. Cinematic tone, Leica photo quality, ultra-detailed textures.”
Figure 8: An example image output from SD 2.1 using the Japanese café prompt.
Prompt Specificity Experiments.
To evaluate the effect of prompt specificity on fingerprinting for SD models, we use the following prompt regimes:
  • Diverse Prompts: Randomly sampled from DiffusionDB (Wang et al. 2023) on HuggingFace (poloclub/diffusiondb).
  • Category-Specific Prompts: “People” category prompts from PartiPrompts (Yu et al. 2022) on HuggingFace (nateraw/parti-prompts).
  • Focused Subset: A small set of generic human-related prompts: “a child,” “a man,” “a woman,” “a girl,” “a person,” “a boy”.
  • Generic Single Prompt: “a high quality photo”.
  • Specific Single Prompt: The Japanese café prompt described above.
Generation Configuration.
By default, SD images are generated using 25 inference (denoising) steps, classifier-free guidance scale of 7.5, and fp16 precision.
Training Details.
Reconstructors are trained using the Adam optimizer to minimize MSE between predicted (reconstructed) and ground-truth pixel values at secret indices. Learning rates are set to $10^{-3}$ for StyleGAN2 experiments and $10^{-4}$ for SD experiments for best convergence.
PGD Setups in Forgery Attacks.
We evaluate the forgery threat on a reconstructor trained for a target StyleGAN2 model (FFHQ, 70k, ADA) at fingerprint length 32, against four negative models: (i) 1k with ADA, (ii) 30k with ADA, (iii) 70k with BCR, and (iv) 70k without augmentation. As shown in Figure 2 (left), all negative variants yield 0% FPR@95%TPR at fingerprint length 32.
PGD attacks are configured with an $\ell_{\infty}$ perturbation budget of $\epsilon=0.5$, a step size of 0.001, and momentum of 0.9. Attacks are run for up to 500 steps with early stopping upon successful evasion. These settings are used across all forgery experiments unless otherwise specified. We evaluated multiple PGD configurations by sweeping over different levels of step sizes and momentum values, and this setup yielded the strongest attack performance against baseline detectors. All configurations consistently result in a 0% attack success rate against the AuthPrint detector.
Hardware.
All experiments are conducted on NVIDIA H200 GPUs.
Appendix B Impact of Denoising Steps on Detection Performance in SD
As shown in Figure 9, reducing the number of denoising steps significantly reduces detection error for SD models.
Figure 9: Impact of the number of denoising steps on the performance of AuthPrint for SD models. The evaluation uses the Japanese café prompt, with a fingerprint length of 1024, a reconstructor with 674M parameters, and 512k training samples. We compare three inference settings: 25, 15, and 5 steps.
Appendix C Evasion Attacks
The experiment results for the evasion-based forgery attack using PGD are shown in Table 1.
@@ Appendix C Evasion Attacks
Table 1: PGD-based evasion attack results under model substitution. Each group reports: success rate (SR, %), average PGD steps, and LPIPS distance for successfully attacked images. “–” indicates no successful attack within 500 steps. The positive model is FFHQ-70k (ADA aug.).
Negative Case Baseline Yu-2019 AuthPrint
SR Steps LPIPS SR Steps LPIPS SR Steps LPIPS
@@
Appendix D Fingerprint Recovery Attacks
Let $x\in\mathbb{R}^{d}$ denote an image sampled from the certified generator, and $x^{\prime}\in\mathbb{R}^{d}$ a manipulated version of $x$. Let $S\subseteq[d]$ be the secret fingerprint index set of size $l:=|S|$, and $M\subseteq[d]$ be the set of manipulated pixel indices. Let $R:\mathbb{R}^{d}\to\mathbb{R}^{l}$ denote the fingerprint reconstructor, with prediction $R(x)=r\in\mathbb{R}^{l}$. Let $\tau$ be the needed increase in MSE for a rejection to take place, and define $\delta^{2}:=\mathbb{E}[(x_{i}-x^{\prime}_{i})^{2}]$ as the expected pixel perturbation variance over manipulated indices.
Derivation of Rejection Threshold.
Here, we derive the rejection threshold kk—defined as the minimum number of pixel indices in the manipulated set $M$ that must overlap with the true fingerprint set $S$ (i.e. the minimum $|M\cap S|$) for the detector to reject a manipulated authentic image.
Under reconstructor stability, $R(x^{\prime})\approx R(x)=r$, thus the change in the MSE before and after manipulation is:

### v2

@@ IV Experimental Setup
[…]
Detection error. We quantify the performance of AuthPrint in detecting the audited model by measuring detection error as the FPR at 95% TPR, treating samples of the targeted model as instances of the positive class and samples of the other models as instances of the negative class. For simplicity, we report detection error results for a single run throughout the paper, as the variance across random seeds was consistently below 1%.
Generation and Training Setup. SD images are generated with 25 denoising steps, classifier-free guidance scale 7.5, and fp16 precision. Reconstructors are trained with Adam optimizers to minimize MSE between predicted and ground-truth fingerprint values. Learning rates are set to $10^{-3}$ for StyleGAN2 and $10^{-4}$ for SD.
Prompt Specificity Study. To study the role of prompt conditioning in guided models, we evaluate 5 prompt regimes for SD spanning different levels of prompt specificity:
  1. I.
    Diverse prompts from DiffusionDB [25],
  2. II.
    “People”-category prompts from PartiPrompts [26],
  3. III.
  4. IV.
    A single fixed generic prompt (“a high quality photo”),
  5. V.
    A single fixed specific prompt (Japanese café)
These regimes progressively constrain the output distribution, allowing us to assess how prompt specificity impacts detection.
Forgery Attacks. To evaluate evasion-based forgery, we apply projected gradient descent (PGD) attacks [27] against a detector whose reconstructor is trained on a StyleGAN2 target model (FFHQ-70k-ADA) with fingerprint length 32. Negative classes are derived from four model variants: 1k-ADA, 30k-ADA, 70k-BCR, and 70k-no-aug. The PGD attack is configured with an $\ell_{\infty}$ budget of $\epsilon=0.5$, step size 0.001, momentum 0.9, and a maximum of 500 iterations, with early stopping once a successful forgery is found. Since we assume a threat model in which the detector is hidden from the adversary, we simulate the attack using a CNN classifier (with ResNet-18 structure [28]) trained to discriminate between each negative distribution and the target distribution. The gradient of this surrogate ResNet-18 is then used to guide PGD in perturbing images so as to deceive the true detector. We compare this detector’s robustness to two baseline detectors, described in the corresponding results section.
Software and Hardware. All experiments are implemented in PyTorch and executed on NVIDIA H200 GPUs. Training an AuthPrint pipeline on a StyleGAN2 model typically requires 4-6 GB GPU memory, while for SD it requires 17-23 GB GPU memory.
Reproducibility. Following open science best practices, we will publish a repository with the documentation and code for all our experiments upon acceptance of the paper. The repository will include our implementation of AuthPrint, data processing helper methods, and the datasets used. We will also provide detailed instructions on how to reproduce the results of our experiments.
V Results
This section provides a comprehensive empirical evaluation of AuthPrint, organized around key design choices, scalability, and robustness. We first analyze the role of fingerprint length in detection accuracy across GAN- and diffusion-based generators under model substitution. We then examine the effects of scaling training resources—specifically reconstructor capacity and dataset size—on performance, and evaluate ensemble detection as a means of mitigating false positives. Next, we assess AuthPrint under model modifications such as pruning and quantization, and investigate the influence of prompt specificity in conditional generative models like SD. Finally, we study adversarial robustness, focusing on forgery attacks instantiated via evasion and fingerprint recovery.
Our findings reveal several consistent trends. First, appropriate tuning of fingerprint length is critical: intermediate lengths achieve the best balance between signal strength and reconstructability, yielding consistently low detection error (e.g. 0 for StyleGAN2 and near-0 for SD). Second, detection performance on diffusion models improves markedly with larger reconstructors and increased training samples, underscoring the importance of scalability in narrowing the gap with GANs. Third, ensemble detection and prompt specificity both contribute to stronger performance. Finally, AuthPrint demonstrates resilience to adaptive adversaries under realistic threat models, with security deriving from the secrecy of fingerprint indices and the inherent difficulty of reconstructing ordered fingerprints under a black-box setting. Collectively, these results establish AuthPrint as a practical and reliable method for model authentication across diverse generative architectures and adversarial conditions.
V-A Effect of Fingerprint Lengths
To investigate the effect of fingerprint length on AuthPrint’s detection performance, we first conduct an ablation study by measuring AuthPrint’s detection error across varying fingerprint lengths, while holding all other factors constant. We conduct this ablation study for both families of models: GAN-based (StyleGAN2) and diffusion-based (SD) models (see Section IV for details).
Figure 2 shows the detection error for StyleGAN2 models, evaluated on FFHQ (left) and LSUN-Cat (right). Fingerprint lengths increase in powers of 2 to cover the model’s output space. StyleGAN2 target and negative models share the same architecture but differ in training configurations such as dataset size and data augmentation. Thus, this setup probes AuthPrint’s sensitivity to subtle changes in the training procedure that induce visually imperceptible variations in image quality.
Figure 3 shows the results of the evaluation for SD models, where the target model is SD 2.1, and negatives are earlier versions (SD 1.5 to 1.1), introducing progressively larger distributional shifts due to architectural and dataset differences. The maximum fingerprint size is larger compared to StyleGAN2 due to the larger output size of SD models.
Figure 2: AuthPrint detection performance on model substitution for pre-trained StyleGAN2 models (left: FFHQ, right: LSUN-Cat) across varying fingerprint lengths (number of secret pixels chosen). Target and negative models share the architecture but differ in training data size or augmentation, as detailed in the figure. The $x$-axis is in log-scale.
As seen in the figures, AuthPrint consistently achieves low detection error across both model families, demonstrating robustness in separating target from non-target outputs. Detection performance peaks at intermediate fingerprint lengths: 16–32 for StyleGAN2 and 64 for SD. Very short fingerprints provide too little signal for the reconstructor to reliably recover the fingerprint, while excessively long fingerprints create redundancy that dilutes useful information and reduces accuracy under fixed model capacity. These results suggest that there is a sweet spot in fingerprint length, where the signal is rich enough to capture model-specific dependencies but compact enough to be reconstructed effectively. In practice, this implies that the verifier can tune fingerprint length to find an optimal accuracy.
Figure 3: AuthPrint detection performance on model substitution for SD models across varying fingerprint lengths (number of secret pixels chosen). The target is SD 2.1; negative models are earlier versions (SD 1.5, 1.4, 1.3, 1.2, 1.1) with lower image qualities. The Japanese café prompt is used throughout. The $x$-axis is in log-scale.
[…]
@@ V-A Effect of Fingerprint Lengths
Another factor contributing to the performance gap is the generative process itself. GANs often leave subtle artifacts or statistical fingerprints that AuthPrint may implicitly exploit [29]. Diffusion models, by contrast, progressively suppress such artifacts through iterative denoising during the generation process. Figure 4 illustrates this effect by comparing detection results under different inference schedules—25, 15, and 5 denoising steps—where each “step” corresponds to one iteration of the reverse diffusion process during image generation. The AuthPrint reconstructor is trained on SD 2.1 as the target model and evaluated against five negative models (SD 1.1–1.5). Across all negatives, reducing the number of denoising steps consistently lowers detection error, suggesting that shallower sampling may retain more model-specific traces for SD models that aid detection.
Figure 4: Impact of the number of denoising steps (i.e. iterations of reverse diffusion) on the performance of AuthPrint for SD models. The evaluation uses the Japanese café prompt, with a fingerprint length of 1024, a reconstructor with 674M parameters, and 512k training samples. We compare three inference settings: 25, 15, and 5 steps.
V-B Scaling Up Training Enhances Detection
@@
As shown in Figure 5, we evaluated reconstructors with 32M, 187M, and 674M parameters on SD models (left subfigure), holding all other variables constant. The smallest (32M) failed to reject negative samples, indicating insufficient capacity for the fingerprinting task. In contrast, performance improved substantially with larger reconstructors, and the 674M model achieved near-perfect detection error on SD 1.1.
We also varied the reconstructor’s training set size (128k, 320k, 512k samples) and observed vanishing detection error with more data (right subfigure). These results show that scaling both reconstructor capacity and training volume significantly improves performance. In our most demanding setting (674M reconstructor, 512k samples), training took $\sim$240 GPU hours on an NVIDIA H200—costly for academic clusters but feasible in industrial or governmental settings. Importantly, training is only needed once per target model.
@@
V-C Boosting Performance with Ensemble Detection
To reduce FPR, AuthPrint can be naturally extended with an ensemble detection scheme, where multiple reconstructors are independently trained using different fingerprint index vectors. At verification, each image is evaluated by all reconstructor-index pairs, and the final detection error is defined as the worst-case MSE across the ensemble, i.e., $e(x^{\prime})=\max_{i\in[N]}\text{MSE}(R_{\phi_{i}}(x^{\prime}),(x^{\prime})_{s_{i}})$, where each $R_{\phi_{i}}$ is trained on $s_{i}$, and $N$ is the ensemble size. This conservative strategy boosts sensitivity: while a negative sample may yield low error on one reconstructor-index pair, it is unlikely to do so across all pairs, thereby reducing false acceptance rates.
@@ V-C Boosting Performance with Ensemble Detection
We evaluate this scheme on SD models using identical hyperparameters but distinct index sets per reconstructor. As shown in Figure 6, detection error decreases with ensemble size, with diminishing returns suggesting convergence. The ensemble design is flexible and can be extended to heterogeneous configurations with varying fingerprint lengths, architectures, or training setups. We leave exploration of the parameters of ensemble approaches to future work.
V-D Detection of Target Model Modifications
@@
Figure 7 shows the relationship between detection error (FPR) and image quality (FID) under all optimizations. AuthPrint achieves effective trade-offs under pruning and downsampling. For example, moderate pruning (62.5%) increases FID modestly to 14.69 but drives FPR to near-zero. Similarly, downsampling from $\times256$ to $\times224$ sharply reduces FPR despite only slight increases in FID. Quantization, by contrast, appears more robust to detection: high precision loss (INT4) renders high FID (>> 200) and leads to 0 FPR as expected, but moderate quantization (INT6) results in high FPR despite quality degradation (FID $\approx$ 53). These findings suggest that while AuthPrint is highly effective at detecting distribution shifts induced by pruning and downsampling, quantization presents a greater challenge. With appropriate tuning of fingerprint length and other pipeline hyperparameters, or through ensemble strategies, deployment in real-world settings may be made even more robust against model compression techniques. This opens a promising avenue for future investigations.
V-E Prompt Specificity Enhances Detection for SD
@@
This raises a key question for AuthPrint: How does prompt specificity affect detection performance? We evaluate detection error under five prompt regimes, ordered by increasing specificity (see Section IV for prompt details):
@@
Figure 8: Effect of prompt specificity on detection performance for SD. We compare five regimes: I. Diverse prompts from DiffusionDB; II. Category-specific “People” prompts from PartiPrompts; III. A focused subset of II (see Section IV); IV. A single generic prompt (“a high quality photo”); V. A single specific prompt (Japanese café prompt).
@@
These results follow a pattern: diverse prompts lead to a broad, overlapping output space, making it harder for the detector to distinguish models. Specific prompts constrain the output distribution, making model differences more detectable. Importantly, a high performance only for specific prompts does not limit practical applicability. Many real-world applications—e.g., profile portraits, product photos, architectural renderings, or avatars—rely on consistent, repeatable prompts, where our approach can be effective.
V-F Evaluation of Adaptive Adversarial Attacks
V-F1 Evasion Attacks
A common strategy for bypassing detectors is evasion, typically implemented via gradient-guided perturbations such as Projected Gradient Descent (PGD) [27]. In our context, an evasion attack corresponds to forgery: an adversary perturbs an image sampled from a negative (non-target) generator so that the detector classifies it as coming from the target model.
We evaluate this threat in a model-substitution scenario. The AuthPrint detector’s reconstructor is trained on a StyleGAN2 target model (FFHQ-70k-ADA) with fingerprint length 32. Negative samples are drawn from four substitute StyleGAN2 variants (1k-ADA, 30k-ADA, 70k-BCR, and 70k-no-aug), which—when the reconstructor is trained on the target—each achieve 0% FPR at 95% TPR (see Figure 2, left). In this threat model we assume the adversary has full access to the target and negative generators but only black-box interaction with the detector itself.
@@ V-F1 Evasion Attacks
    Baseline: A ResNet-18 classifier trained directly to distinguish target outputs from negatives (no fingerprinting).
@@
    Yu et al. (2019): A state-of-the-art, public, output-independent model-fingerprinting method [15].
@@
    AuthPrint: Our method, which uses a reconstructor trained with secret, output-dependent fingerprints.
Because the detector is assumed inaccessible, the adversary performs a surrogate-based, black-box attack: they train a surrogate detector and use its gradients to guide PGD on inputs sampled from the negative generators. Concretely, for each of the three detector designs we train a surrogate ResNet-18 classifier to distinguish the corresponding negative distribution from the target distribution, then run PGD using the surrogate’s gradients to craft perturbations. PGD hyperparameters are given in Section IV.
For each detector we report three quantities: (i) the attack success rate (fraction of negative images that the attack converts to be accepted as target), (ii) the average number of PGD steps to success, and (iii) the LPIPS perceptual distance between original and forged images. These metrics jointly capture how effective the forgery is and how perceptible the required perturbation is.
@@
TABLE I: PGD-based evasion attack results under model substitution. Three detectors are evaluated: (i) Baseline: a baseline detector (ResNet-18), (ii) Yu-2019: a model fingerprinting-based detector proposed by Yu et al. [15], (iii) AuthPrint: our method. Each group reports: attack success rate (ASR, %), average PGD steps, and LPIPS distance for successfully attacked images. “–” indicates no successful attack within 500 steps. The positive model is FFHQ-70k-ADA.
Negative Model Baseline Yu-2019 AuthPrint
ASR Steps LPIPS ASR Steps LPIPS ASR Steps LPIPS
@@
Results show that AuthPrint is resistant to this attack: the attacker fails to generate even a single successful forgery. As shown in Table I, both baseline detectors are defeated by PGD, with a 100% attack success rate—i.e., all non-authentic images are misclassified as authentic. Moreover, the attacks maintain LPIPS consistently below $0.01$ for the baseline detectors, meaning that the attacks are visually imperceptible. While the Yu-2019 detector requires slightly more steps to evade, both are ultimately ineffective. In contrast, AuthPrint achieves a 0% attack success rate. The attacker’s inability to access the secret pixel indices or the reconstructor’s parameters prevents them from training a viable surrogate. This secrecy is vital for the security of AuthPrint: if the attacker had white-box access to the reconstructor, or knowledge of the fingerprint index vector, a PGD attack would likely succeed.
V-F2 Fingerprint Recovery Attacks
Beyond evasion-based attacks, an adversary might also seek to recover the fingerprint indices $s$ in order to forge outputs. We decompose this attack into two steps: first, recovering the fingerprint set $S:=\mathrm{set}(s)$; second, searching for the ordered vector $s$ from $S$. Both the set and the order are crucial—without any one of them, the adversary cannot replicate the behavior of $R$.
Step 1: Recover the set $S$. To recover the set $S$, one can probe the detector through pixel manipulation: starting from an authentic image $x\in\mathbb{R}^{d}$, the attacker manipulates a subset $M\subset[d]$ of $m$ pixels (e.g., setting them to zero) to produce a modified image $x^{\prime}$, and observes whether it passes or is rejected by the verifier. The detector of AuthPrint is designed to compute $D(x;R,s)=\mathrm{MSE}[R(x),x_{s}]$, which remains low upon manipulation if two conditions hold: (1) the reconstructor output $R(x)$ is stable to the manipulation, and (2) the manipulated pixels $M$ do not significantly overlap with the fingerprint set $S$.
1.1: Reconstructor Stability Assumption. Throughout the analysis, we assume that $R(x)$ is stable under small manipulations. To validate this empirically, we generate 100 authentic images using a StyleGAN2 model (FFHQ-70k-ADA), and apply 1,000 random manipulation trials per image. For each trial, we randomly select $m\in\{1,2,4,\ldots,65536\}$ pixel indices to modify, and compute the mean $\ell_{1}$ distance between the reconstructor outputs $R(x)$ and $R(x^{\prime})$. Results in Figure 9 confirm that reconstructor outputs remain stable for small values of $m$.
[…]
1.2: Detector Rejection Criterion. We then formalize the detector’s rejection criterion upon pixel manipulation by analyzing the change in MSE between $x$ and $x^{\prime}$. Let $l=|S|$ be the fingerprint length, and let $\delta^{2}:=\mathbb{E}[(x_{i}-x^{\prime}_{i})^{2}]$ denote the expected pixel perturbation variance over manipulated indices. Under reconstructor stability assumption (i.e., $R(x^{\prime})\approx R(x)=r$), the change in MSE due to pixel manipulation is:
$\Delta\mathrm{MSE}(x^{\prime},x)=\frac{1}{l}\sum_{i\in S}\left[(r_{i}-x^{\prime}_{i})^{2}-(r_{i}-x_{i})^{2}\right].$ (1)
Since only manipulated fingerprint pixels affect the MSE, we can simplify Eq. 1: